Clean Company is a little short of cash so they are planning to do an 8% stock dividend instead. Callie has 300 shares of Clean and they are currently selling at $55.00 per share. Clean has a total of 20,000 shares outstanding.

Before the stock dividend, what is the total value of Callie’s investment in Clean Company? ___________________

After the stock dividend, Callie will have ______________ shares.

After the stock dividend, the shares will be worth ___________ per share.

After the stock dividend, the total value of Callie’s investment in Clean Company will be _____________________.
Answer #1

Shares after stock dividend = current shares*(1+stock dividend%) = 300*(1+0.08) = 324

Price after stock dividend = current price/(1+stock dividend%) = 55/(1+0.08) =50.925

Value = price*shares = 50.925*324 = 16500